Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%?
Please help me on this

Answer by Nate(3500)

P*R*T = I
T = 1
P*R = I
I + I = 2720
x*0.14 + (20,000 - x)*0.13 = 2720
0.14x + 2600 - 0.13x = 2720
0.01x = 120
x = $12,000 in 14%